uzsbgck:

498.82 centigrams to milligrams ( use scientific notation)

hulkster123:

4998.2 milligrams

uzsbgck:

Ok but what about the interger for x 10

uzsbgck:

@hulkster123

AZ:

 What will the integer be?

uzsbgck:

3

AZ:

Remember: when we're moving to the left, our integer will be positive If we were moving towards the right, the integer for the exponent would have to be negative

AZ:

So what's your final answer?

uzsbgck:

4998.2 x 10 ^ -3

AZ:

not quite we moved that decimal point 3 places to the left so it would be 4.9882 (it's also 4.9882 not 4.9982, that one user who commented earlier made a typo) But does that make sense? We're moving the decimal point sometimes to the left, other times to the right And in doing so we're multiplying some number by a 10^something

uzsbgck:

wait so the 3 is postiive ? I get the first part

Nevaeh12:

use this site is can help you https://www.mathpapa.com/simplify-calculator/

AZ:

yes because we're moving to the left another way to think about it what is 10^3? It's 1000 If we multiply 1 by 1000, you'll get 1000 So when you move the decimal to the left (example 1000.0 to 1.0 *10^3), you're making that number smaller because you're pulling out 1000 from it and writing it as 10^3

AZ:

And similarly, it works the other way if you have a negative exponent like -2 10^(-2) = 1/10^2 = 1/100 We're essentially DIVIDING by 100 So that means the number was already small to begin with so if we have 0.03 we want to move that decimal point over to the RIGHT to get 3.0 we would have to multiply by 10^(-2) so 0.03 = 3 * 10^(-2)

AZ:

Do those examples help clarify it for you? @uzsbgck